The earth rotates once per day about an axis passing through the north and south poles, an axis that is perpendicular to the plane of the equator. Assuming the earth is a sphere with a radius of 6.38 x 106 m, determine the speed and centripetal acceleration of a person situated (a) at the equator and (b) at a latitude of 57.0 ° north of the equator.
Answer #1

a)

a = r w^2 = 6.38* 10^6* ( 2 * 3.14 / 24* 60* 60)^2

a = 0.0337 m/s^2

======

b)

radius from the axis of rotation

r = 6.38 cos 30 * 10^6 = 5.525* 10^6

now

a = 5.525* 10^6* ( 2* 3.14 / ( 24 * 3600)) ^2

a = 0.0292 m/s^2

======

Comment before rate in case any doubt, will reply for sure.. goodluck